OpenStudy (anonymous):

can someone show me how to do this??...find the sum of the given arithmetic series 30+23+16+9+...+(-33) a. -15 b. 315 c. -3 d. -30

OpenStudy (anonymous):

Here a = 30, d = 23 - 30 = -7 Use: \[\large a_n = a + (n-1)d\] So: \[30 + (n - 1)(-7) = -33 \implies -7n = -70\] \[n = 10\] Now use: \[\large Sum = \frac{n}{2}(a + a_n)\]

OpenStudy (anonymous):

notice first: 30+23+16+9+...+(-33) -33 +...+ 9+16+23+30 ----------------------- 3 +3 +...........+3 + 3 then find how many terms you series has multiply 3 by this number and devide by 2

OpenStudy (anonymous):

\(a_n\) is the last term here that is -33.. n = 10 here.. a = 30 here..

OpenStudy (anonymous):

oo okay i see, my formula was wrong! i was divided a+an by 2 instead of the n!

OpenStudy (anonymous):

thanks for the help:))

OpenStudy (anonymous):

It is okay..